OpenStudy (anonymous):

Find the x-, y, and z-intercepts for the equation. 2x - 6y - 3z = 12

OpenStudy (neonumbrella5115):

What do you think the answer is? What have you tried so far?

OpenStudy (anonymous):

(6, -2, -4) This... But it wasn't right... I don't think I'm doing it right.

OpenStudy (neonumbrella5115):

So you need to figure out when x and y are 0, then what is x. As a first step

OpenStudy (anonymous):

Is it 6?

OpenStudy (neonumbrella5115):

No, try again. -3z=12

OpenStudy (tkhunny):

This, my friends, is why we acquaint ourselves with the "Intercept Form". 2x - 6y - 3z = 12 Divide by 12 \(\dfrac{x}{6}+\dfrac{y}{-2}+\dfrac{z}{-4} = 1\) Done. The three intercepts are now staring at you.

OpenStudy (anonymous):

(6, -2, -4) I had the answer right originally I guess. Just was putting parenthesis... The answer was actually right.

OpenStudy (tkhunny):

Do NOT forget the intercept form.

OpenStudy (anonymous):

Huh? o__o

OpenStudy (anonymous):

Find the x-, y-, and z-intercepts for the equation. Enter your answers separated by commas; example: -2, 3.5, 4 3x – 5y + 10z = 15 I have another. How would I do this in intercept form??

OpenStudy (tkhunny):

Maybe divide by 15 and make sure all the operations are "+"? Go forth!! Divide!

OpenStudy (anonymous):

3x, 5y, and 10z? I divide them all by 15 right?

OpenStudy (anonymous):

0.2, 0.3, 0.6 ? O_o

OpenStudy (tkhunny):

...and make sure there is only addition. Look at my previous example. I had to alter some subtraction.

OpenStudy (anonymous):

Erm... I don't get it I am sorry

OpenStudy (tkhunny):

Let's see your division.

OpenStudy (anonymous):

I used a calculator to divide the numbers by 15 and got those 0. numbers above

OpenStudy (tkhunny):

Why would you use a calculator? That's just silly. Do the arithmetic. Are you allergic to fractions?

OpenStudy (anonymous):

it's okay I am going to skip it for now and come back to it on Monday when my teacher is available D: I'm more of a visual learner, I don't understand math very well. Sorry if I sound like an idiot... I can't even divide in my head, and yeah... Fractions confuse the heck out of me..

OpenStudy (tkhunny):

Oh, come on! Really? Not even going to try it? Oh, well. Don't forget to mention the "Intercept Form" to your teacher.

OpenStudy (anonymous):

LOL I did try... I just don't want to be made to feel stupid if I don't know how to do something

OpenStudy (anonymous):

I will mention it. Thank you for helping me.

OpenStudy (tkhunny):

No "make you feel stupid" going on around here. Are you sure? Gotta' get up to speed on the fractions. They are NOT your enemy. Honest. :-)